# CliffordNet: All You Need is Geometric Algebra

### v1

@@ 4.1 Main Results and Efficiency Analysis
Table 1 presents the comparison with state-of-the-art efficient backbones. Our CliffordNet-Base achieves 78.05%, establishing a new SOTA for models under 4M parameters, surpassing ResNet-18 (11.2M) by a clear margin.
The Cost of Geometry vs. Depth. We observe an interesting computational trade-off. Transitioning from Nano (2 shifts) to Fast (5 shifts) improves accuracy by $\sim$1.4% but increases training time by $\sim$56%. In contrast, adding an FFN layer (Fast $\to$ Base) only adds $\sim$11% overhead. This indicates that the current bottleneck lies in the geometric sampling density (number of shifts).
Potential for Acceleration. It is crucial to note that our current implementation of the “Shifted Geometric Product” relies on standard PyTorch primitives (e.g., torch.roll and slicing), which are memory-bound operations involving redundant data movement. Unlike standard convolutions which benefit from decades of cuDNN optimization, our geometric primitives are unoptimized. This high latency for multi-shift interactions suggests a significant room for improvement. By implementing custom fused CUDA/Triton kernels to perform rolling interactions on-the-fly in registers (avoiding VRAM round-trips), we expect the cost of increasing shifts to become negligible, potentially doubling the throughput of the Base/Fast models in future iterations.
@@
Table 2: Comparison with Efficient Backbones on CIFAR-100. All models are trained from scratch using the same rigorous recipe (200 epochs, AutoAugment, WD=0.05). CliffordNet outperforms strong baselines across all parameter scales.
@@ 2.5M Scale (Tiny)
CliffordNet-Fast 2.6M 0.0 77.63%
>>3.0M Scale (Base)
@@
CliffordNet-Base 3.0M 1.0 78.05%
Analysis. As shown in Table 2, while ShuffleNetV2 achieves a strong accuracy of 74.60% through its ingenious engineering heuristics, our CliffordNet-Nano surpasses it by 1.81% with comparable parameters. This result suggests that the implicit channel mixing induced by our mathematically grounded Clifford interaction (via the shifted geometric product) is more effective than the explicit, handcrafted channel shuffling used in traditional efficient CNNs. Our Fast model (Ratio=1) matches the performance of ResNet-18 while using only 24% of the parameters. Even the No-FFN variant achieves 77.63% accuracy, surpassing MobileNetV2 by a large margin, validating our hypothesis that geometric interactions can replace generic MLPs. To prove that our efficiency stems from the architecture and not the dataset, we trained a standard ViT-Tiny with ‘mlp_ratio=1.0‘. The model collapsed to 65.87% accuracy (Table 2), whereas CliffordNet maintained 78+%. This confirms that standard attention lacks the geometric density to function without heavy FFNs.
@@
Table 3: Ablation on Interaction Density and FFN Necessity. We compare variants of CliffordNet with different geometric shifts and MLP ratios. Increasing geometric shifts (2 $\to$ 5) yields significant gains (+1.22%), whereas adding an explicit FFN only provides marginal improvement (+0.42%), confirming that the geometric interaction layer is the primary driver of representation learning.
Model Variant FFN (Ratio) Shifts Params Top-1 Acc Interpretation
Baseline (MobileNetV2) - - 2.3M 70.90% Reference
CliffordNet-Nano No (0.0) 2 ([1,2]) 1.43M 76.41% High Efficiency
CliffordNet-Fast No (0.0) 5 ([1..15]) 2.61M 77.63% Optimal Balance
CliffordNet-Base Yes (1.0) 5 ([1..15]) 3.01M 78.05% Diminishing Returns
Interaction Density vs. FFN Capacity. To disentangle the contributions of geometric interaction density (number of shifts) and explicit channel mixing (FFN), we analyze three variants of CliffordNet in Table 3.
First, we observe that even with minimal geometric priors ($\mathcal{S}=\{1,2\}$), CliffordNet-Nano achieves 76.41%, outperforming the MobileNetV2 baseline (70.90%) by 5.51% while using $\sim$38% fewer parameters (1.43M vs. 2.3M). This underscores the inherent efficiency of the Clifford interaction layer: mere local geometric reasoning captures more discriminative features than standard depthwise convolutions.
Second, increasing the density of geometric sampling proves highly effective. Expanding the shift set from 2 to 5 (creating CliffordNet-Fast) boosts accuracy by +1.22% (76.41% $\to$ 77.63%). This significant gain confirms that capturing long-range channel dependencies (e.g., $s=15$) via the rolling mechanism enriches the feature representation substantially.
Most notably, comparing “Fast” (No FFN) with “Base” (Ratio=1) reveals a phenomenon of diminishing returns. Reintroducing the FFN layer adds $\sim$0.4M parameters but yields a negligible gain of only 0.42% (77.63% $\to$ 78.05%). This empirically validates our hypothesis that the Clifford Interaction layer — equipped with multiplicative wedge products and non-linear gating — already possesses sufficient representational capacity, rendering the explicit FFN largely redundant.
Table 4: Component Analysis: Scalar Energy vs. Bivector Structure. We disentangle the Clifford Interaction into Dot-only and Wedge-only variants (both  1.7M params). The Dot product captures feature magnitude (energy) via diagonal terms ($u_{i}u_{i}$), while the Wedge product is strictly anti-symmetric ($u_{i}\wedge u_{i}=0$), relying solely on structural relations. Remarkably, the Wedge variant rivals the Dot variant despite lacking explicit energy information, and their combination (CliffordNet) yields the best performance.
@@
Dot-Only Symmetric (Has Energy) Coherence 76.91%
Wedge-Only Anti-Sym (No Energy) Structure 76.35%
CliffordNet Geometric Complete 77.63%
Inner Product vs. Exterior Product. To disentangle the distinct contributions of scalar and bivector components, we compare Dot-Only and Wedge-Only variants under identical settings ($\mathcal{S}=\{1,2,4,8,15\}$, 1.68M params).
Table 4 shows that Dot-Only (76.91%) slightly outperforms Wedge-Only (76.35%). However, this comparison reveals a profound geometric insight regarding Diagonal Energy:
@@
    Dot Product (Energy + Relation): The generalized inner product includes near-diagonal correlations. Crucially, it implicitly encodes feature magnitude or energy (akin to $x_{i}^{2}$), which is fundamental for activation-based deep learning.
@@
    Wedge Product (Pure Relation): The exterior product is strictly anti-symmetric ($x\wedge x=0$). It inherently discards all self-magnitude (energy) information and relies solely on cross-channel geometric relations.
Remarkable Robustness of Bivectors. Despite operating without explicit energy information — effectively fighting with “one hand tied behind its back” — the Wedge-Only variant achieves performance nearly on par with the Dot-Only baseline. This suggests that the structural topology (captured by bivectors) is almost as discriminative as the feature intensity itself. The combination of the two (CliffordNet) thus provides the complete picture: Energy (Dot) + Structure (Wedge).
Context Receptive Field & Depth. We analyze the instantiation of the local geometric context $\mathcal{C}(H)$ and the impact of the context stream’s capacity. A baseline using a single $3\times 3$ depth-wise convolution achieves 77.05%. Expanding the receptive field with a large $7\times 7$ kernel improves performance to 77.28%, confirming the importance of capturing macroscopic geometric structures. However, our proposed factorized design (stacking two $3\times 3$ kernels) yields the best performance at 77.63%. This indicates that the depth of the diffusion process (iterative refinement) is more critical than the instantaneous receptive field size. This factorization provides a dual benefit: It reduces the parameter count of the context stream. It acts as an iterative diffusion process, which we empirically find to be more robust than a single large-step update. Consequently, we adopt the factorized $3\times 3$ stack as our default design.

### v2

@@ 4.1 Main Results and Efficiency Analysis
Table 1 presents the comparison with state-of-the-art efficient backbones. Our CliffordNet-Lite achieves 79.05%, establishing a new SOTA for models under 3M parameters, surpassing ResNet-18 (11.2M) by a clear margin.
The Cost of Geometry vs. Depth. We observe an interesting computational trade-off. Transitioning from Nano (2 shifts) to Lite (5 shifts) improves accuracy by $\sim$1.2% but increases training time by $\sim$56%. This indicates that the current bottleneck lies in the geometric sampling density (number of shifts). However, with CUDA acceleration applied to the Clifford Interaction, the training times for both the Nano and Lite variants become comparable (57.64 mins vs. 58.27 mins).
@@
Table 2: Comparison with Efficient Backbones on CIFAR-100. CliffordNet outperforms strong baselines across all parameter scales.
@@ 2.5M Scale (Tiny)
CliffordNet-Lite 2.6M 0.0 79.05%
>>3.0M Scale (Small)
@@
ResNet-50 23.7M - 79.14%
DenseNet-121 7.0M - 80.20%
CliffordNet-32(3 shifts,cli_mode=’full’) 4.8M 0.0 81.42%
CliffordNet-64(5 shifts,cli_mode=’inner’) 8.6M 0.0 82.46%
Analysis. As shown in Table 2, while ShuffleNetV2 achieves a strong accuracy of 74.60% through its ingenious engineering heuristics, our CliffordNet-Nano surpasses it by 4.3% with comparable parameters. This result suggests that the implicit channel mixing induced by our mathematically grounded Clifford interaction (via the shifted geometric product) is more effective than the explicit, handcrafted channel shuffling used in traditional efficient CNNs. Our small models match or outperform the performance of ResNet-18 and MobileNetV2 by a large margin, validating our hypothesis that geometric interactions can replace generic MLPs. To prove that our efficiency stems from the architecture and not the dataset, we trained a standard ViT-Tiny with ‘mlp_ratio=1.0‘. The model collapsed to 65.87% accuracy (Table 2), whereas CliffordNet-Lite maintained 79+%. This confirms that standard attention lacks the geometric density to function without heavy FFNs.
[…]
All the models here are trained from scratch, without using any pre-trained weights or transfer learning techniques, and without using warm-up training epochs.
@@
Table 3: Ablation on Interaction Density and gFFN Necessity. We compare variants of CliffordNet with different geometric shifts, context modes and gFFNs.
Model Variant Shifts Params Top-1 Acc
Baseline (MobileNetV2) - 2.3M 70.90%
CliffordNet-Nano(abs) 2 ([1,2]) 1.43M 76.41%
CliffordNet-Lite(abs) 5 ([1..16]) 2.61M 77.63%
CliffordNet-Nano(diff) 2 ([1,2]) 1.43M 77.82%
CliffordNet-Lite(diff) 5 ([1..16]) 2.61M 79.05%
CliffordNet-Nano(diff)+gFFN-G 2 ([1,2]) 2.22M 78.22%
CliffordNet-Lite(diff)+gFFN-G 5 ([1..16]) 3.40M 79.57%
Impact of Context Mode and Interaction Density. To disentangle the contributions of geometric interaction density (number of shifts) and interaction modes, we analyze the variants of CliffordNet. Table 3 reveals two key insights. First, the Differential Mode (pure $\Delta H$) consistently outperforms the Absolute Mode by $\sim$1.4%, confirming that representation learning benefits from filtering out static self-energy. Second, increasing the geometric sampling density (Shifts $2\to 5$) and integrating the Global Geometric FFN (+gFFN-G) yields further gains. Notably, even our minimal Nano(diff) variant achieves 77.82%, surpassing the MobileNetV2 baseline by nearly 7% with 38% fewer parameters.
Table 4: Component Analysis: Scalar Energy vs. Bivector Structure. We disentangle the Clifford Interaction into Inner-only and Wedge-only variants (both  1.63M params). The Inner product captures feature magnitude (energy) via diagonal terms ($u_{i}u_{i}$), while the Wedge product is strictly anti-symmetric ($u_{i}\wedge u_{i}=0$), relying solely on structural relations. Remarkably, the Wedge variant rivals the Inner variant despite lacking explicit energy information, and their combination (CliffordNet) yields the best performance.
@@
Inner-Only Symmetric (Has Energy) Coherence 78.17%
Wedge-Only Anti-Sym (No Energy) Structure 77.76%
CliffordNet Geometric Complete 79.05%
Inner Product vs. Exterior Product. To disentangle the distinct contributions of scalar and bivector components, we compare Inner-Only and Wedge-Only variants under identical settings ($\mathcal{S}=\{1,2,4,8,16\}$). All reported results here utilize the Differential Mode.
Table 4 shows that Inner-Only (78.17%) slightly outperforms Wedge-Only (77.76%). However, this comparison reveals a profound geometric insight regarding Diagonal Energy:
@@
    Inner Product: The generalized inner product includes near-diagonal correlations. Crucially, it implicitly encodes feature magnitude or energy (akin to $x_{i}^{2}$), which is fundamental for activation-based deep learning.
@@
    Wedge Product: The exterior product is strictly anti-symmetric ($x\wedge x=0$). It inherently discards all self-magnitude (energy) information and relies solely on cross-channel geometric relations.
Remarkable Robustness of Bivectors. Despite operating without explicit energy information, the Wedge-Only variant achieves performance nearly on par with the Inner-Only baseline. This suggests that the structural topology (captured by bivectors) is almost as discriminative as the feature intensity itself. The combination of the two (CliffordNet) thus provides the complete picture: Energy (Inner) + Structure (Wedge).
Data-Driven Geometric Selection. By providing the full spectrum of geometric interactions (both scalar coherence and bivector structure), CliffordNet avoids hard-coding inductive biases regarding which feature type is more discriminative. Instead, the learnable projection layer $\mathcal{P}$ acts as a soft selector, dynamically weighting the contribution of the inner and exterior products based on the data distribution. For texture-heavy tasks, the network may prioritize the wedge product; for shape-biased tasks, it may favor the inner product. This algebraic completeness ensures that the optimal geometric representation is always within the hypothesis space of the model, maximizing generalization across diverse visual domains. For resource-constrained deployment, structural pruning could be applied to selectively remove either the scalar or bivector pathways if the target task relies predominantly on one geometric mode.
@@ 6 Conclusion and Future Work
  • Optimized Geometric Topology (Decoupled Sampling): Currently, we share the same shift set $\mathcal{S}$ for both inner and exterior products. However, our ablation study suggests a degree of informational redundancy between the scalar and bivector streams. A promising direction is to decouple the interaction topologies ($\mathcal{S}_{dot}\neq\mathcal{S}_{wedge}$). By assigning distinct receptive fields to the coherent (Inner) and structural (Wedge) pathways — essentially sampling independent diagonals of the interaction matrix — we could maximize the information entropy of the layer without increasing computational cost.